sunlight:

The monthly salary of a person id $12,000 and his monthly expenditure is $8,500. Find the ratio if his: i) salary to expenditure. ii) Expenditure to savings iii) Savings to salary

KynxKatz:

Solution: Monthly salary of a person = Rs. 12000 Monthly expenditure = 8500 Saving of the person = 12000 - 8500 = Rs. 3500 (i) Ratio between salary and expenditure 12000:8500= 12000/8500= 120/85=24/17=24:17 (ii) Ratio between expenditure and savings 8500:3500=85/35=17/7= 17:7 (iii) Ratio between savings and salary 3500:12000= 3500/12000= 35/120=7/24= 7:24

KynxKatz:

Number one, the monthly salary for person is 12,000 and is monthly expenditure is rupees 3500 find the ratio of his salary to expenditure. Now. First of all, the salary is given to us as 12,000. Okay, and is monthly expenditure is given to us as 8005 so we have to find this ratio because this is a salary first one is a salary and this next one is a spend. It's beautiful. Sure. So the ratio will come out to be which for let's cancel out the Z goes from both places. Let's divide both of them five by five. So by Phi will give you 24 and 85 by 5 will give you 10 and 7 so really we 70 so this is your final answer is 24 is 217 for the first one because you're the first one here answer is 24 is 270. Let's go to the second question. Please note it down. This is the second the second question is expenditures to a savings. So first of all, see events a salary - expenditure so 12,000 - 8,500 so 12,000 - 8500 RS series. So this will give you answer to be 3500. So he sees total 3500 in a month. Now we have to find the ratio of his expenditure to a savings. So is expenditure is 8,500 and savings is 3 this find so the ratio will be 8500 to 3,500. So first of all, let's cancel out the Rose this will give you 17 when divided by 5. And this will give you seven. So your answer to this. This one is 17 is 2/7. Okay. Now, let's move on to the last one that is the ratio of savings to salary. So saving the game is 3500 and his salary is 12,000 Civic find the ratio. So first of all, let's cancel out the zeros and this is directly answer because we cannot further divide both of the new police or 35 is 212 is the final answer the Thunder. https://www.lidolearning.com/questions/m-bb-selina6-ch11-ex11b-q1/q1-the-monthly-salary-of-a-per/